OpenStudy (anonymous):

If f(x) = 4 sin(x) + 2x^x, find f′( 2 ).

OpenStudy (phi):

I assume you can take the derivative of sin(x) for x^x write x as \[ x= e^{\ln(x)} \] and \[ x^x= \left(e^{\ln(x)}\right)^x =e^{x\ln(x)} \]

OpenStudy (anonymous):

so the derivative is 4 cos(x) + 2(e^(xln(x)))?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

4 cos(x) is ok \[\frac{d}{dx} e^u = e^u \frac{d}{dx}u \] I don't see the the derivative of x ln(x)

OpenStudy (anonymous):

@phi so it is 4cosx+ e^(xln(x))*(1+ln(x)?

OpenStudy (phi):

yes. but remember that to find the derivative we used \[ x^x= e^{x\ln(x)} \] so you could rewrite the answer as \[ 4 \cos x + x^x(1+\ln(x)) \] notice the question says " find f′( 2 ). " so you need to replace x with 2 and simplify

OpenStudy (phi):

also, the original equation f(x) = 4 sin(x) + 2x^x has a 2 * x^2 you are missing the 2 out front of the 2nd term

OpenStudy (phi):

evaluate at x=2 \[ 4 \cos x + 2x^x(1+\ln(x)) \]